# Emergent Weyl-like points in periodically modulated systems and Floquet engineering

### v2

@@ Abstract
We investigate a three-dimensional topological phase resembling a Weyl semimetal, modulated by a periodic potential and engineered through Floquet dynamics. This system is constructed by stacking two-dimensional Chern insulators and hosts Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. The Weyl-semimetal-like phase exhibits characteristics akin to those of Weyl semimetals, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the presence of Fermi arcs connecting the Weyl-like points. Unlike traditional Weyl semimetals, these features manifest in real space rather than momentum space. Additionally, we compute the layer-resolved Floquet Hall conductivity and demonstrate that the positions of the Weyl-like points can be controlled via high-frequency laser pumping.
@@ I Introduction
Stacks of two-dimensional (2D) Chern insulator phases have proven to be a versatile platform for realizing extraordinary topological phases, attracting considerable attention in recent years Mogi et al. (2017a, b); Xiao et al. (2018); Mogi et al. (2022); Qin et al. (2023); Chen and Shen (2024); Chang et al. (2013); Yoshimi et al. (2015); Shan et al. (2010); Lu et al. (2010); Li et al. (2010); Lu et al. (2013); Chen et al. (2019); Sun et al. (2020); Liu et al. (2010); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Sun et al. (2023); Qin et al. (2022a); Zhu et al. (2023). For instance, the well-known Weyl semimetal phase can be interpreted as a stack of 2D Chern insulators with orbital-dependent interlayer tunneling Yang et al. (2011); Chen et al. (2015, 2018a, 2018b); Li et al. (2021). Similarly, phases such as the axion insulator and the hinged quantum spin Hall effect arise from alternating stacks of Chern insulators with opposite Chern numbers Chen et al. (2021, 2024, 2023a, 2023b). Moreover, higher-Chern-number quantum anomalous Hall insulators can be constructed by alternating stacks of Chern insulators and normal insulators Zhao et al. (2022); Baba et al. (2022); Li et al. (2022, 2023).
Weyl semimetals are characterized by linearly dispersing band-touching points, known as Weyl points, in momentum space Wan et al. (2011); Chan et al. (2016); Yan and Felser (2017); Guo et al. (2023). These Weyl points are topologically protected and act as monopoles or antimonopoles of the Berry curvature Li et al. (2016); Lu and Shen (2017). According to the no-go theorem Nielsen and Ninomiya (1981a, b), Weyl points always appear in pairs and are connected by surface states known as Fermi arcs.
Recent advancements have realized ideal Weyl semimetals across various platforms. They have been implemented in electrical circuits Lu et al. (2019); Lee et al. (2020), ultracold atomic gases with three-dimensional (3D) spin-orbit coupling Dubček et al. (2015); Roy et al. (2018); Wang et al. (2021); Lu et al. (2020), photonic crystals Lu et al. (2015); Li et al. (2018), and acoustic metamaterials Xiao et al. (2015); Yang and Zhang (2016); Peri et al. (2019).
Zhang et al. recently proposed and studied a 2D phase modulated by a periodic potential Zhang et al. (2023). This system exhibits edge modes with spectral flow dependent on position, in contrast to traditional Chern insulators, where the spectral flow of edge modes depends on momentum. Moreover, they suggested that the periodic potential could be experimentally realized through the effects of charge density waves Huang et al. (2021); Qin et al. (2020).
@@
In this work, we propose and study a 3D topological phase resembling a Weyl semimetal, modulated by a periodic potential and realized through a simple stack of 2D Chern insulators. The system features Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. This Weyl-semimetal-like phase exhibits key properties of a Weyl semimetal, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arcs connecting these points. However, these behaviors are observed in real space, rather than momentum space, as in traditional Weyl semimetals. Additionally, we calculate the layer-resolved Floquet Hall conductivity under high-frequency periodic driving and show that the positions of the Weyl-semimetal-like nodes can be tuned by high-frequency laser driving.
The structure of this paper is as follows: In Section II, we introduce the Weyl-like lattice Hamiltonian with a periodic potential along the $z$-direction, along with the conventional Weyl semimetal model, and discuss the duality between these two models. In Section III, we derive the Floquet effective Hamiltonians for both models under high-frequency laser pumping. Section IV presents calculations of the Floquet Weyl nodes, the layer Floquet Chern number, the Floquet energy spectrum, and the Floquet Fermi arc for both models under high-frequency laser driving. A brief discussion on potential experimental realizations of the Weyl-like model is provided in Section V. Finally, Section VI offers a summary of the key results.
@@ II Model
We begin with the two-band model Hamiltonian describing a stack of 2D Chern insulators in a periodic potential along the $z$-direction,
$H\!=\!\sum_{j_{z}}\left[h_{\text{2D}}({\bf k}_{||})\!+\!V(z)\sigma_{z}\right]c_{j_{z}}^{{\dagger}}c_{j_{z}},$ (1)
where $c_{j_{z}}^{{\dagger}}$ is the electron creation operator in the $j_{z}$-th layer, $h_{\text{2D}}({\bf k}_{||})\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Chern insulator Hamiltonian when $m_{z}\times m_{0}<0$, and ${\bf k}_{||}\!=\!(k_{x},k_{y})$. Here, $a$ is the lattice constant, and $\sigma_{i}$ $(i\!=\!x,y,z)$ are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z=j_{z}a$ is the position along the $z$-direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the periodic wavelength. In the following calculations, we use the fixed parameters $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm. The detailed derivation of the matrix form for the real-space model in a periodic potential along the $z$-direction, with open boundary conditions (OBCs) in $y$-direction and periodic boundary conditions (PBCs) in $x$-direction, is provided in Section SI in the Supplemental Material Sup .
Figure 1: Duality between the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column). (a1) Schematic representation of the decoupled Weyl-like model in real space, featuring a spatially periodic potential $V(z)=V_{z}\cos(Q_{z}z)$, where $Q_{z}=2\pi/\lambda_{z}$. (b1) Weyl-like cones in $(k_{x},k_{y},z)$ space at two critical planes, $(k_{x},k_{y},z_{c-})$ and $(k_{x},k_{y},z_{c+})$, marking topological transitions. (c1) Layer-resolved Chern number as a function of layer $z$ for the decoupled Weyl-like model with $V_{z}=1$ eV, $\lambda_{z}=L_{z}=31a$, and the lattice constant $a=1$ nm. (d1) Energy spectrum (blue solid curves) for the Weyl-like model with $k_{x}=k_{y}=0$. (e1) Energy spectrum (blue solid curves) and Fermi-like arc (blue horizontal line for $E=0$) for the Weyl-like model under OBCs along the $y$-direction, with $k_{x}=0$ and system size $L_{y}=31a$. (a2) Schematic representation of the Weyl semimetal model in real space, with a momentum-periodic potential $U(k_{z})=t_{z}\cos(k_{z}a)$. (b2) Weyl cones in the $(k_{x},k_{y},k_{z})$ momentum space at two critical planes $(k_{x},k_{y},-k_{c})$ and $(k_{x},k_{y},+k_{c})$, indicating the topological features of the model. (c2) Chern number as a function of momentum $k_{z}$ for the Weyl semimetal model with $t_{z}=1$ eV. (d2) Energy spectrum (black solid curves) for the Weyl semimetal model with $k_{x}=k_{y}=0$. (e2) Energy spectrum (black solid curves) and Fermi arc (black horizontal line for $E=0$) for the Weyl semimetal model under OBCs along the $y$-direction, with $k_{x}=0$ and system size $L_{y}=31a$. The other parameters are $m_{z}=0.5$ eV, $m_{0}=1$ eV, $t_{x}=t_{y}=t_{0}=1$ eV, and $L_{z}=31a$.
@@ II.1 Weyl-like model
The spectral flow of the system is described by the Weyl-like model Hamiltonian density, Eq. (1), defined in the $(k_{x},k_{y},z)$-space as follows:
@@
$\!+t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}.$
As shown in Fig. 1(a1), the system can be interpreted as an array of isolated 2D systems, each with a different topological mass: $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)$, where $V(z)\!=\!V_{z}\cos(Q_{z}z)$.
To study the topological properties of the system, we define the layer-resolved Chern number for the $m$-th occupied band as Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
@@
where the non-Abelian Berry curvature is given by Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
@@
where ${\cal A}_{k_{\alpha}}^{mn}({\bf k}_{||},z)$ is the Berry connection of the occupied bands with $m,n\in{\rm occ}$, defined as
@@
Here, “occ” is used to denote the occupied bands. A detailed derivation of the non-Abelian Berry curvature in (4) is provided in Section SII in the Supplemental Material Sup .
The layer-resolved Chern number is illustrated in Fig. 1(c1). Figures 1(a1) and 1(b1) show that each 2D system behaves as a Chern insulator with a Chern number $|C|=1$ for $z_{c-}<z<z_{c+}$, and as a normal insulator with $C=0$ for $z<z_{c-}$ and $z>z_{c+}$. The critical values of $z$, denoted by $z_{c\pm}$, are determined by the condition $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)\!=\!0$ and are given by:
$\frac{z_{c\pm}}{\lambda_{z}}\!=\!\frac{1}{2\pi}\left[\pi\pm\arccos\left(\frac{m_{z}}{V_{z}}\right)\right],$ (6)
where the condition $-1\leqslant\tilde{m}_{z}/V_{z}\leqslant 1$ is satisfied.
Figure 1(d1) shows the energy spectrum as a function of position $z$ with $k_{x}\!=\!k_{y}\!=\!0$. Notably, a band inversion is observed near $z\!=\!z_{c\pm}$, clearly illustrating the spectral flow of the stacked system. Additionally, the edge states (zero-energy modes) can be observed in Fig. 1(e1) that are dual to Fig. 1(e2).
II.2 Weyl semimetal
To facilitate a comparative study, we review the Weyl semimetal model Hamiltonian, which takes the following form Yang et al. (2011); Chen et al. (2015, 2018a, 2018b)

### v3

@@ Abstract
We investigate a three-dimensional (3D) topological phase resembling a Weyl semimetal, modulated by a periodic potential and engineered through Floquet dynamics. This system is constructed by stacking two-dimensional Chern insulators and hosts Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. The Weyl-semimetal-like phase exhibits characteristics akin to those of Weyl semimetals, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the presence of Fermi arcs connecting the Weyl-like points. Unlike traditional Weyl semimetals, these features manifest in real space rather than momentum space. Furthermore, we calculate the local density of states, the layer Hall conductance, and the total 3D Hall conductivity, demonstrating that the Weyl-semimetal-like phase remains stable under weak and moderate interlayer couplings. Finally, we compute the Floquet Chern number and show that the positions of the Weyl-like points can be tuned via high-frequency laser pumping.
@@ I Introduction
Stacks of two-dimensional (2D) Chern insulator phases have proven to be a versatile platform for realizing extraordinary topological phases, attracting considerable attention in recent years Mogi et al. (2017a, b); Xiao et al. (2018); Mogi et al. (2022); Qin et al. (2023); Chen and Shen (2024); Chang et al. (2013); Yoshimi et al. (2015); Shan et al. (2010); Lu et al. (2010); Li et al. (2010); Lu et al. (2013); Chen et al. (2019); Sun et al. (2020); Liu et al. (2010); Dabiri et al. (2021); Dabiri and Cheraghchi (2021); Sun et al. (2023); Qin et al. (2022a); Zhu et al. (2023). For instance, the well-known Weyl semimetal phase can be interpreted as a stack of 2D Chern insulators with orbital-dependent interlayer tunneling Yang et al. (2011); Chen et al. (2015, 2018a, 2018b); Li et al. (2021). Similarly, phases such as the axion insulator and the hinged quantum spin Hall effect arise from alternating stacks of Chern insulators with opposite Chern numbers Chen et al. (2021, 2024a, 2023a, 2023b). Moreover, higher-Chern-number quantum anomalous Hall insulators can be constructed by alternating stacks of Chern insulators and normal insulators Zhao et al. (2022); Baba et al. (2022); Li et al. (2022, 2023).
Weyl semimetals are characterized by linearly dispersing band-touching points, known as Weyl points, in momentum space Wan et al. (2011); Chan et al. (2016); Yan and Felser (2017); Guo et al. (2023). These Weyl points are topologically protected and act as monopoles or antimonopoles of the Berry curvature Li et al. (2016); Lu and Shen (2017). According to the no-go theorem Nielsen and Ninomiya (1981a, b), Weyl points always appear in pairs and are connected by surface states known as Fermi arcs. Recent advancements have realized ideal Weyl semimetals across various platforms. They have been implemented in electrical circuits Lu et al. (2019); Lee et al. (2020), ultracold atomic gases with three-dimensional (3D) spin-orbit coupling Dubček et al. (2015); Roy et al. (2018); Wang et al. (2021); Lu et al. (2020), photonic crystals Lu et al. (2015); Li et al. (2018), and acoustic metamaterials Xiao et al. (2015); Yang and Zhang (2016); Peri et al. (2019).
Zhang et al. recently proposed and studied a 2D phase modulated by a periodic potential Zhang et al. (2023). Unlike conventional Chern insulators, where the spectral flow of edge modes is momentum-dependent, their system exhibits edge modes whose spectral flow depends on spatial position. They also suggested that such a periodic potential could be experimentally realized via charge density wave effects Huang et al. (2021); Qin et al. (2020). In a related development, Chen et al. explored a 2D higher-order topological insulator and demonstrated that the energies of the corner modes can be tuned by a periodic spin-orbital potential Chen et al. (2024b).
@@
In this work, we propose and study a 3D topological phase resembling a Weyl semimetal, modulated by a periodic potential and realized through a simple stack of 2D Chern insulators. The system features Weyl-like points defined in the parameter space $(k_{x},k_{y},z)$, distinct from the conventional Weyl points in momentum space $(k_{x},k_{y},k_{z})$. This Weyl-semimetal-like phase exhibits key properties of a Weyl semimetal, including linear dispersion near the Weyl-like points, nontrivial bulk topology, and the emergence of Fermi arc connecting these points. However, these behaviors are observed in real space, rather than momentum space, as in traditional Weyl semimetals. We further compute the local density of states, the layer Hall conductance, and the total 3D Hall conductivity, demonstrating that the Weyl-semimetal-like phase remains robust under weak and moderate interlayer couplings. Additionally, we evaluate the Floquet Chern number under high-frequency periodic driving and show that the positions of the Weyl-like nodes can be effectively tuned by high-frequency laser fields.
The structure of this paper is as follows. In Section II, we introduce the Weyl-like model Hamiltonian with a periodic potential along the $z$-direction, alongside the conventional Weyl semimetal model, and discuss the duality between these two models. Section III presents calculations of the local density of states, the layer Hall conductances, and the total 3D Hall conductivity in the presence of interlayer hopping for the Weyl-like model. In Section IV, we derive the Floquet effective Hamiltonians for both models in Section II under high-frequency laser driving. Section V details the analysis of Floquet Weyl nodes, Floquet Chern numbers, the Floquet energy spectra, and Floquet Fermi arcs for both models. A brief discussion of possible experimental realizations of the Weyl-like model is given in Section VI. Finally, Section VII summarizes the main results of the paper.
@@ II Model
In this section, we will introduce the Weyl-like model and the Weyl semimetal model.
@@ II.1 Weyl-like model
We begin by introducing a two-band model Hamiltonian that describes a stack of 2D Chern insulators subjected to a periodic potential along the $z$-direction. This model is defined in the $(k_{x},k_{y},z)$ space as
@@
where $h_{\text{2D}}({\bf k}_{||})\!=\!m_{z}\sigma_{z}\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]\sigma_{z}\!+\!t_{x}\sin(k_{x}a)\sigma_{x}\!+\!t_{y}\sin(k_{y}a)\sigma_{y}$ represents the Hamiltonian of a 2D Chern insulator when $m_{z}\times m_{0}<0$ Shen (2017), and ${\bf k}_{||}\!=\!(k_{x},k_{y})$. Here, $a$ is the lattice constant, and $\sigma_{i}$ $(i\!=\!x,y,z)$ are Pauli matrices acting on the orbital space. The spatially periodic potential is given by $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $V_{z}$ is the potential strength, $z\!=\!j_{z}a$ denotes the position of the $j_{z}$-th layer along the $z$-direction, and $\lambda_{z}\!=\!2\pi/Q_{z}$ is the wavelength of the modulation.
We emphasize that the layers are decoupled in this case, ensuring that $z$ remains a good quantum number. The effects of interlayer coupling will be addressed in Section III.
Figure 1: Duality between the decoupled Weyl-like model (left column) and the Weyl semimetal model (right column). The two green vertical dashed lines indicate the positions of the two Weyl-like (or Weyl) points. (a1) Schematic of the decoupled Weyl-like model in real space, featuring a spatially periodic potential $V(z)\!=\!V_{z}\cos(Q_{z}z)$, where $Q_{z}\!=\!2\pi/\lambda_{z}$. (b1) Weyl-like cones in the $(k_{x},k_{y},z)$ space at two critical planes, $(k_{x},k_{y},z_{c-})$ and $(k_{x},k_{y},z_{c+})$, marking topological transitions. (c1) Chern number of each layer [Eq. (3)] as a function of layer index $z$ for the decoupled Weyl-like model with $V_{z}\!=\!1$ eV, $\lambda_{z}\!=\!L_{z}\!=\!31a$, and the lattice constant $a\!=\!1$ nm. (d1) Energy spectrum [Eq. (6)] (blue solid curves) of the Weyl-like model at $k_{x}\!=\!k_{y}\!=\!0$. (e1) Energy spectrum (blue solid curves) and Fermi-like arc (blue horizontal line for $E\!=\!0$) for the Weyl-like model under open boundary conditions (OBCs) along the $y$-direction, with $k_{x}\!=\!0$ and system size $L_{y}\!=\!31a$. (a2) Schematic of the Weyl semimetal model in real space, featuring a momentum-dependent potential $U(k_{z})\!=\!t_{z}\cos(k_{z}a)$. (b2) Weyl cones in momentum space $(k_{x},k_{y},k_{z})$ at two critical planes $(k_{x},k_{y},-k_{c})$ and $(k_{x},k_{y},+k_{c})$, indicating the topological features of the model. (c2) Chern number [Eq. (II.2)] as a function of $k_{z}$ for the Weyl semimetal model with $t_{z}\!=\!1$ eV. (d2) Energy spectrum [Eq. (13)] (black solid curves) of the Weyl semimetal model at $k_{x}\!=\!k_{y}\!=\!0$. (e2) Energy spectrum (black solid curves) and Fermi arc (black horizontal line at $E\!=\!0$) for the Weyl semimetal model under OBCs along the $y$-direction, with $k_{x}\!=\!0$ and system size $L_{y}\!=\!31a$. Other parameters used: $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, and $L_{z}\!=\!31a$.
As illustrated in Fig. 1(a1), the system described by the Weyl-like Hamiltonian in Eq. (1) can be interpreted as an array of decoupled 2D layers, each with a spatially dependent effective mass term $\tilde{m}_{z}(z)\!=\!m_{z}\!+\!V(z)$. For the calculations presented in this work, we adopt the fixed parameters: $m_{z}\!=\!0.5$ eV, $m_{0}\!=\!1$ eV, $t_{x}\!=\!t_{y}\!=\!t_{0}\!=\!1$ eV, $V_{z}\!=\!1$ eV, and $\lambda_{z}\!=\!31a$ with $a\!=\!1$ nm. The detailed matrix representation of the real-space Hamiltonian with the periodic potential along the $z$-direction – assuming open boundary conditions (OBCs) along the $y$-direction and periodic boundary conditions (PBCs) along the $x$-direction – is provided in Section SI of the Supplemental Material Sup .
The eigenenergies of the Weyl-like model described by Eq. (1) are given by
$E_{\pm}\!=\!\pm\sqrt{[t_{x}\sin(k_{x}a)]^{2}\!+\![t_{y}\sin(k_{y}a)]^{2}\!+\![\tilde{M}_{z}(k_{x},k_{y},z)]^{2}},$
where $\tilde{M}_{z}(k_{x},k_{y},z)\!=\!\tilde{m}_{z}(z)\!+\!m_{0}[2\!-\!\cos(k_{x}a)\!-\!\cos(k_{y}a)]$.
Since the layers are decoupled from each other, we can define a $z$-dependent Chern number of each layer for the $m$-th occupied band as Fukui et al. (2005); Sun et al. (2020); Otrokov et al. (2019); Gao et al. (2021); Chen et al. (2024a); Mong et al. (2010); Essin et al. (2009); Wang et al. (2015); Varnava and Vanderbilt (2018); Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
@@
where the non-Abelian Berry curvature is defined as Chen and Shen (2024); Fu et al. (2021); Price et al. (2015); Zhang et al. (2019)
@@
and the Berry connection ${\cal A}_{k_{\alpha}}^{mn}({\bf k}_{||},z)$ of the occupied bands (with $m,n\in{\rm occ}$) is given by
@@
Here, “occ” denotes the set of occupied bands. A detailed derivation of the non-Abelian Berry curvature in Eq. (4) is provided in Section SII in the Supplemental Material Sup .
The $z$-dependent Chern number of each layer is illustrated in Fig. 1(c1). Figures 1(a1) and 1(b1) show that each layer behaves as a Chern insulator with Chern number $|C|\!=\!1$ for $z_{c-}<z<z_{c+}$, and as a normal insulator with $C\!=\!0$ outside this region (i.e., for $z<z_{c-}$ and $z>z_{c+}$). To determine the critical values $z_{c\pm}$, we evaluate the eigenenergies of the Weyl-like Hamiltonian (1) at the high-symmetry point $k_{x}\!=\!k_{y}\!=\!0$, obtaining from Eq. (II.1)
$E_{\pm}\!=\!\pm\tilde{m}_{z}(z)\!=\!\pm\left[m_{z}\!+\!V_{z}\cos(Q_{z}z)\right].$ (6)
Furthermore, the critical values $z_{c\pm}$ correspond to the gap-closing points where $E_{\pm}\!=\!\pm\tilde{m}_{z}(z)\!=\!\pm\left[m_{z}\!+\!V_{z}\cos(Q_{z}z)\right]\!=\!0$, leading to
$z_{c\pm}\!=\!\frac{\lambda_{z}}{2\pi}\left[\pi\pm\arccos\left(\frac{m_{z}}{V_{z}}\right)\right],$ (7)
under the condition $-1\leqslant m_{z}/V_{z}\leqslant 1$.
Figure 1(d1) shows the energy spectrum as a function of position $z$ at $k_{x}\!=\!k_{y}\!=\!0$. A clear band inversion occurs near $z\!=\!z_{c\pm}$, highlighting the spectral flow characteristic of the layered topological structure. Additionally, Fig. 1(e1) reveals the presence of a Fermi arc, further demonstrating the system’s topological nature.
Moreover, the total 3D Hall conductivity for the $m$-th occupied band is given by Burkov and Balents (2011)
[…]
where $\sigma_{H}^{2D}(z)\!=\!(e^{2}/h)C_{m}(z)$ is the Hall conductance of each layer associated with the $m$-th occupied band, $C_{m}(z)$ is defined in Eq. (3), and $L_{z}$ denotes the system size along the $z$-direction. Using the critical values $z_{c\pm}$ obtained in Eq. (7), we can simplify the 3D Hall conductivity expression as
[…]
where the approximation in the second line assumes $L_{z}\gg a$, allowing the sum to be replaced by an integral over $z$. This result highlights how the 3D Hall response is governed by the range of layers exhibiting nontrivial topology.
II.2 Weyl semimetals
To facilitate a comparative study, we now review the Weyl semimetal model, whose Hamiltonian is given by Yang et al. (2011); Chen et al. (2015, 2018a, 2018b)